1)form a pde by eliminating the arbitrary function from z=exp(2x+3y)f(2x-3y)


1
Expert's answer
2021-05-31T12:09:32-0400

We multiply by "exp(-(2x+3y))" and receive: "z\cdot exp(-(2x+3y))=f(2x+3y)". After denoting "w=z\cdot exp(-(2x+3y))" we receive: "w_x=2f'" , "w_y=3f'". We receive the equation: "3w_x=2w_y", "w_x=z_x\cdot exp(-(2x+3y))-2z\cdot exp(-(2x+3y))" and "w_y=z_y\cdot exp(-(2x+3y))-3z\cdot exp(-(2x+3y))". Finally, we receive: "3z_x\cdot exp(-(2x+3y))-6z\cdot exp(-(2x+3y))=2z_y\cdot exp(-(2x+3y))-6z\cdot exp(-(2x+3y))"

Thus, finally, we get the equation: "3z_x=2z_y"